The equation y=-1777x+27153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is, x=0 corresponds to 2000, x=2 corresponds to 2002, x=5 corresponds to 2005, and so on. Predict the number of gun deaths in 2005 and 2008. In what year will the number of gun deaths be 11,160?
Answer by stanbon(75887)

The equation y=-1777x+27153 can be used to predict the number y of gun deaths in the united states x years after 2000, that is, x=0 corresponds to 2000, x=2 corresponds to 2002, x=5 corresponds to 2005, and so on. Predict the number of gun deaths in 2005 and 2008. In what year will the number of gun deaths be 11,160?
----------------------
d(x)=-1777x+27153
d(5) = -1777*5 + 27153 = 18268
---
d(8) = -1777*8 + 27153 = 12937
--------
Solve 11,160 = -1777x + 27153
-1777x = -15993
x = 9
--
Year = 2009
================================
Cheers,
Stan H.